Fermi National Laboratory
 
Operation Summary for November 21, 2011 - November 23, 2011
 

Monday November 21

The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the MiniBooNE, NuMI, and the Meson FTBFT-978 experiments taking beam.

During the day shift, at 8:56 AM, a Linac Klystron vernier solenoid power supply tripped off; it reset. At 9:30 AM, Linac personnel decided they had to change the vernier solenoid power supply, which held off beam for about 1.5 hours. Operators prepared NuMI for an absorber room access to check their water skid. A MiniBooNE person asked Operations to turn off their beam in preparation for an access. Operations resumed sending beam to NuMI at 11:02 AM. Operators reported problems with the MI-8 beam at 3:08 PM; MI and operators investigated.

During the evening shift, Operations established beam to MiniBooNE at 4:02 PM. Operators reported at 5:19 PM that the problems with the MI-8 line disappeared before anyone could discover the cause. At 6:37 PM, a Booster RF station (BRF1) tripped off for the first of two times in five minutes; it reset. A Booster person investigated. Operations disabled the T-978 beam at 7:03 PM. Operators rebooted the Linac Low Level RF system at 7:11 PM.

Tuesday November 22

The midnight shift began with Operations sending beam to the NuMI and MiniBooNE.

During the midnight shift, at 2:26 AM, the NuMI beam permit tripped off on a BPM (E:65BPMB); operators rebooted the BPM. Operation established beam to T-978 at 6:57 AM. Operators reported at 7:57 AM that a NuMI BPM (NBPM6N) stopped polling; operators rebooted the BPM.

During the day shift, a Muon person (formerly the Antiproton Source) began a study period around 9:31 AM. RF personnel began work on BRF1 at 9:48 AM. Operations reported at 11:34 AM that MTA had taken beam this morning. A Switchyard person changed the T-978 beam from 16 GeV to 12 GeV at 3:18 PM. The Muon study ended at 3:28 PM.

During the evening shift, Operations disabled beam to T-978 at 7:43 PM. At 7:43 PM, the I- Source gate valves closed and the chopper tripped off; operators investigated and found that the I- Source vacuum gauge was off. Operators turned it on and everything reset. Operators completed a MI tune at 8:08 PM. Operators completed a 750 KeV line tune at 10:09 PM. Operators reported short downtimes due to LRF1, LRF2, and KRF4.

Wednesday November 23

The midnight shift began with Operations sending beam to the NuMI and MiniBooNE.

During the midnight shift, at 5:05 AM, the NuMI beam permit tripped off when the output current of a power supply (I:HP608N) drifted high; everything reset. At 7:06 AM, operators reported that a Booster CAMC crate 25 had stopped responding; Controls personnel investigated and replaced the crate’s power supply. Booster resumed sending beam at 7:29 AM.

Operations Coordinator (OpsCo) Plans for Today and the Holiday Weekend

The OpsCo approved of the plans by Booster, MI, and Muon personnel to conduct studies.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/23/2011    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory